### v1

# On the equivalence of NMDS codes

Abstract
An $[n,k,d]$ linear code is said to be maximum distance separable (MDS) or almost maximum distance separable (AMDS) if d=nk+1d=n-k+1 or d=nkd=n-k, respectively. If a code and its dual code are both AMDS, then the code is said to be near maximum distance separable (NMDS). For $k=3$ and $k=4$, there are many constructions of NMDS codes by adding some suitable projective points to arcs in $\mathrm{PG}(k-1,q)$. In this paper, we consider the monomial equivalence problem for some NMDS codes with the same weight distributions and present new constructions of NMDS codes.
Keywords: NMDS code; Arc; Hyperoval; Monomial equivalence
Mathematics Subject Classification (2020): 94B05 51E21
1 Introduction
An $n,k]_{q}$ linear code $\mathcal{C}$ is a $k$-dimensional subspace of $\mathbb{F}^{n}_{q}$. The parameter $n$ is called the length of $\mathcal{C}$. The vectors in $\mathcal{C}$ are called codewords. Let $c\in\mathcal{C}$ be a codeword. The support of $c=(c_{1},c_{2},\dots,c_{n})$ is defined as $supp(c)=\{i\in[n]:c_{i}\neq 0\}$, where $[n]$ denotes the set $\{1,2,\dots,n\}$. The weight of $c$ is the cardinality of its support, i.e., $wt(c)=|supp(c)|$. An $n,k,d]_{q}$ linear code $\mathcal{C}$ is an $n,k]_{q}$ linear code with minimum distance $d$, defined as $d=\text{min}_{0\neq c\in\mathcal{C}}wt(c)$. Usually, if the context is clear, the subscript $q$ is omitted by convention in the sequel. A generator matrix of $\mathcal{C}$ is a $k\times n$ matrix $G$ whose rows form a basis of $\mathcal{C}$ as a $\mathbb{F}_{q}$-vector space. The dual code of $\mathcal{C}$ is its orthogonal subspace $\mathcal{C}^{\perp}$ with respect to the Euclidean inner product, defined as $\mathcal{C}^{\perp}=\{x\in\mathbb{F}^{n}_{q}:x\cdot c=0,\forall c\in\mathcal{C}\}$. A monomial matrix over $\mathbb{F}_{q}$ is a square matrix with exactly one nonzero element from $\mathbb{F}_{q}$ in each row and column. Such a matrix $M$ can be expressed as $DP$ or $PD_{1}$, where $D$ and $D_{1}$ are diagonal matrices and $P$ is a permutation matrix. Two $[n,k,d]$ linear codes $C_{1}$ and $C_{2}$, with generator matrices $G_{1}$ and $G_{2}$ respectively, are monomially equivalent if there exist a matrix $L\in\mathrm{GL}(k,q)$ and an $n\times n$ monomial matrix $M$ such that $LG_{1}M=G_{2}$.
The Singleton bound [36] states a relationship among $n,k$ and $d$: $d\leq n-k+1$. The codes meeting this bound are called maximum distance separable (MDS) codes. The famous MDS conjecture states that for a nontrivial MDS code, the maximum length $n$ satisfies: $n\leq q+2$ if $q$ is even and $k=3$ or k=q1k=q-1, and $n\leq q+1$ otherwise [33]. This conjecture was extensively investigated by Ball in [1, 2]. A code with parameters of the form [n,k,nk][n,k,n-k] is said to be almost maximum distance separable (AMDS). If both a linear code and its dual code are AMDS, the code is said to be near maximum distance separable (NMDS), which was introduced by Dodunekov and Landjev [12]. Let $m(k,q)$ denote the maximum possible length for a nontrivial $q$-ary NMDS code. Similarly to the MDS conjecture, the following conjecture for NMDS codes was proposed in [12, 19]: $m(k,q)\leq 2q+2$. Specially, for $k=3$ and $q\geq 8$, it is conjecdtured that $m(3,q)\leq 2q-1$, and it is known m(3,q)=2q1m(3,q)=2q-1 for $q=8,9$.
Both MDS codes and NMDS codes can be investigated in the language of finite geometries. Let $\mathrm{PG}(r,q)$ denote the $r$-dimensional projective space derived from $\mathbb{F}^{r+1}_{q}$, where its points correspond to the one-dimensional subspaces of $\mathbb{F}^{r+1}_{q}$, and its hyperplanes correspond to the $r$-dimensional subspaces of $\mathbb{F}^{r+1}_{q}$. An $n$-arc in $\mathrm{PG}(k-1,q)$ is a set of $n$ points such that every hyperplane intersects the arc in at most k1k-1 points. An arc is said to be maximal if it has the largest possible number of points among all arcs. For an $[n,k,d]$ linear code $\mathcal{C}$ with generator matrix $G$, let $\mathcal{S}_{G}$ be the set of points in $\mathrm{PG}(k-1,q)$ corresponding to the columns of $G$. For any nonzero vector $u\in\mathbb{F}^{k}_{q}$, we can define the hyperplane $\mathcal{H}_{u}=\{x\in\mathbb{F}^{k}_{q}:x\cdot u=0\}$ in $\mathrm{PG}(k-1,q)$. The following lemma establishes a relationship between the weights of codewords in $\mathcal{C}$ and the hyperplanes of $\mathrm{PG}(k-1,q)$.
Lemma 1.1.
[3, Lemma 5.16] Let $u$ be a nonzero vector of $\mathbb{F}^{k}_{q}$. The codeword $uG$ has weight $\omega$ if and only if the hyperplane $\mathcal{H}_{u}$ in $\mathrm{PG}(k-1,q)$ contains $|\mathcal{S}_{G}|-\omega$ points of $\mathcal{S}_{G}$, i.e., $|\mathcal{H}_{u}\cap\mathcal{S}_{G}|=n-wt(uG)$. Especially, any hyperplane in $\mathrm{PG}(k-1,q)$ intersects $\mathcal{S}_{G}$ in at most ndn-d points.
From Lemma 1.1, we see that [n,k,nk+1][n,k,n-k+1] MDS codes and $n$-arcs in $\mathrm{PG}(k-1,q)$ are indeed equivalent objects. Similarly, an [n,k,nk][n,k,n-k] NMDS code can be viewed as a point-set $\mathcal{S}$ of size $n$ in $\mathrm{PG}(k-1,q)$ satisfying the following conditions [12]:
  1. (N1)
    any k1k-1 points of $\mathcal{S}$ generate a hyperplane in $\mathrm{PG}(k-1,q)$;
  2. (N2)
    there exist $k$ points in $\mathcal{S}$ that lie on a hyperplane in $\mathrm{PG}(k-1,q)$;
  3. (N3)
    every $k+1$ points of $\mathcal{S}$ generate $\mathrm{PG}(k-1,q)$.
By adding suitable projective points to the $n$-arcs in $\mathrm{PG}(k-1,q)$, many NMDS codes of dimension $k$ have been constructed. For the case $k=3$, see [40, 20, 21, 42, 13]. For the cases $k=4$ and $k=5$, see [41, 15, 11]. Moreover, there are many NMDS codes with the same parameters derived from these constructions. In this paper, we investigate the monomial equivalence problem for such NMDS codes and provide some new constructions. Specifically, by utilizing hyperovals in $\mathrm{PG}(2,q)$, we propose an infinite family of NMDS codes with length $q+3$ and dimension $3$ for even $q$ (see Theorem 3.9). Through detailed analysis of the action of the homography stabilizers of distinct hyperovals on the points not in hyperovals, we fully determine the equivalence classes of these NMDS codes. By analyzing the orbits of points and planes of $\mathrm{PG}(3,q)$ under the stabilizers of arcs, we provide a unified geometric perspective for constructing several equivalence classes of NMDS codes with length $q+2$ and dimension $4$, derived from arcs in $\mathrm{PG}(3,q)$ with additional points (see Theorems 4.26 and 4.30 ).
This paper is organized as follows. In Section 2, we present some preliminary results on NMDS codes and the maximal arcs in $\mathrm{PG}(2,q)$ and $\mathrm{PG}(3,q)$. The monomial equivalence problem for NMDS codes of dimensions 33 and $4$ is considered in Sections 3 and 4, respectively. In Section 5, we conclude this paper.
2 Preliminaries
In this section, we present some preliminary results which will be used later.
2.1 NMDS codes
Let $\mathcal{C}$ be an [n,k,nk][n,k,n-k] NMDS code with generator matrix $G=(\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{n})$, where $\mathbf{g}_{i}\in\mathbb{F}^{k}_{q}$ and $k\geq 3$. Since $k\geq 3$, the columns of $G$ can be identified with distinct points in $\mathrm{PG}(k-1,q)$. Define $\mathcal{S}_{G}=\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{n}\}$. Then $\mathcal{S}_{G}$ is a point-set in $\mathrm{PG}(k-1,q)$ satisfying the conditions (N1), (N2) and (N3) stated in Section 1.
Let $A_{i}$ be the number of codewords with weight $i$ in $\mathcal{C}$, where $0\leq i\leq n$. The polynomial $A(z)=1+A_{1}z+A_{2}z^{2}+\cdots+A_{n}z^{n}$ is called the weight enumerator of $\mathcal{C}$. The sequence $(1,A_{1},A_{2},\dots,A_{n})$ is termed the weight distribution of $\mathcal{C}$. The following lemma gives the weight distribution formulas for NMDS codes.
Lemma 2.2.
[12] Let $\mathcal{C}$ be an [n,k,nk][n,k,n-k] NMDS code. Then the weight distribution of $\mathcal{C}$ is given by
$A_{n-k+s}=\binom{n}{k-s}\sum_{j=0}^{s-1}(-1)^{j}\binom{n-k+s}{j}\left(q^{s-j}-1\right)+(-1)^{s}\binom{k}{s}A_{n-k}$
for $s\in\{1,2,\ldots,k\}$.
It follows from Lemma 2.2 that the weight distribution of an NMDS code can be determined by the number of its minimum weight codewords, i.e., $A_{n-k}$. Let $\mathcal{H}_{u}$ be the hyperplane in $\mathrm{PG}(k-1,q)$ determined by the nonzero vector $u\in\mathbb{F}^{k}_{q}$. Then by Lemma 1.1, we have
$A_{n-k}=\#\left\{u\in\mathbb{F}_{q}^{k}\setminus\{\mathbf{0}\}:|\mathcal{H}_{u}\cap\mathcal{S}_{G}|=k\right\}.$ (2.1)
Here the notation #A\#A denotes the cardinality of set $A$.
2.2 Hyperovals
An $n$-arc in $\mathrm{PG}(2,q)$ is a set of $n$ points such that no three are collinear. It is know that $n\leq q+2$, with equality holding if and only if $q$ is even. Let $q=2^{m}$ for a positive integer $m$. A hyperoval $\mathcal{HO}$ in $\mathrm{PG}(2,q)$ is an arc with $q+2$ points, where every line in $\mathrm{PG}(2,q)$ intersects $\mathcal{HO}$ in either 0 or $2$ points. Two hyperovals are equivalent if one can be mapped to the other via an automorphism of $\mathrm{PG}(2,q)$. All hyperovals in $\mathrm{PG}(2,q)$ can be constructed using a special type of permutation polynomials over $\mathbb{F}_{q}$, as described in the following Theorem [16].
Theorem 2.3.
Let $q>2$ be a power of $2$. Any hyperoval in $\mathrm{PG}(2,q)$ can be written in the following form
$\mathcal{HO}(f)=\{(1,c,f(c)):c\in\mathbb{F}_{q}\}\cup\{(0,0,1)\}\cup\{(0,1,0)\},$
where $f(x)\in\mathbb{F}_{q}[x]$ is a permutation polynomial of $\mathbb{F}_{q}$ satisfying:
  1. (1)
    $\text{deg}(f)\leq q-2$ and $f(0)=0$, $f(1)=1$;
  2. (2)
    for each $a\in\mathbb{F}_{q}$, $f_{a}(x)=(f(x+a)+f(a))/x$ is also a permutation polynomial of $\mathbb{F}_{q}$, and $f_{a}(0)=0$.
Conversely, every set $\mathcal{HO}(f)$ defined in this way is a hyperoval.
By [6], for even $q$, a conic together with its nucleus (the intersection point of its tangents) forms a hyperoval. Such a hyperoval is said to be regular. As shown by [34], for $q=2,4,8$, every hyperoval is regular. However, for $q=2^{m}$ with $m\geq 4$, irregular hyperovals exist. Any polynomial satisfying the two conditions in Theorem 2.3 is called an oval polynomial (or o-polynomial for short). Table 1 gives a list of all known o-polynomials over $\mathbb{F}_{q}$ with $q=2^{m}$, also see [18, Table 3]. Here,
$F_{1}(x)=\left(\delta^{2}\left(x^{4}+x\right)+\delta^{2}\left(1+\delta+\delta^{2}\right)\left(x^{3}+x^{2}\right)\right)\left(x^{4}+\delta^{2}x^{2}+1\right)^{-1}+x^{1/2};$
$F_{3}(x)=x^{4}+x^{16}+x^{28}+\omega^{11}(x^{6}+x^{10}+x^{14}+x^{18}+x^{22}+x^{26})+\omega^{20}(x^{8}+x^{20})+\omega^{6}(x^{12}+x^{24});$
$F_{2}(x)=\frac{T\left(\beta^{m}\right)(x+1)}{T(\beta)}+\frac{T\left(\left(\beta x+\beta^{q}\right)^{m}\right)}{T(\beta)\left(x+T(\beta)x^{2^{m-1}}+1\right)^{m-1}}+x^{2^{m-1}}$
[…]
Table 1: Hyperovals in $\mathrm{PG}(2,q)$, $q$ even
Name $f(x)$ Conditions Reference
Regular $x^{2}$ [6]
Irregular translation $x^{2^{h}}$ $\gcd(h,m)=1$, $h>1$ [34]
Segre $x^{6}$ $m$ odd [35]
Glynn I $x^{3\times 2^{(m+1)/2}+4}$ $m$ odd [14]
Glynn II $x^{2^{(m+1)/2}+2^{(m+1)/4}}$ $m\equiv 3\pmod{4}$ [14]
$x^{2^{(m+1)/2}+2^{(3m+1)/4}}$ $m\equiv 1\pmod{4}$ [14]
Cherowitzo $x^{2^{e}}+x^{2^{e}+2}+x^{3\times 2^{e}+4}$ $e=(m+1)/2$, $m$ odd [9]
Payne $x^{1/6}+x^{3/6}+x^{5/6}$ $m$ odd [27]
Subiaco $F_{1}(x)$ $\operatorname{Tr}_{q/2}(1/\delta)=1$, $\delta\notin\mathbb{F}_{4}$ if $m\equiv 2\bmod 4$ [30]
Adelaide $F_{2}(x)$ ()(*) [10]
O’Keefe-Penttila $F_{3}(x)$ $\omega^{5}=\omega^{2}+1$, $m=5$ [24]
Some useful properties on hyperovals and o-polynomials are presented below.
Lemma 2.4.
[28, Lemma 4.6.6] If hyperovals $K$ and $K^{\prime}$ in $\mathrm{PG}(2,q)$ have $\frac{1}{2}(q+2)+n$ points in common, with $n>0$, then $K=K^{\prime}$.
Lemma 2.5.
[23] A polynomial $f(x)$ over $\mathbb{F}_{q}$ with $f(0)=0$ is an o-polynomial if and only if $f(x)+ux$ is 2-to-1 for every $u\in\mathbb{F}^{*}_{q}$.
Lemma 2.6.
Let $f(x)\in\mathbb{F}_{q}[x]$ be an o-polynomial, $u_{1},u_{2}\in\mathbb{F}^{*}_{q}$, $a,b\in\mathbb{F}_{q}$ with $(a,b)\neq(0,0)$. Define
$A(u_{1},u_{2})=\#\{x\in\mathbb{F}_{q}:u_{1}f(x)+u_{2}x+u_{1}a+u_{2}b=0\}.$
Then $A(u_{1},u_{2})\in\{0,2\}$ and $\{(u_{1},u_{2})\in(\mathbb{F}_{q}^{*})^{2}:A(u_{1},u_{2})=2\}=(q-1)(q-2)/2.$
Proof.
The proof is similar to that of [42, Lemma 3.1], and so we omit the details. ∎
Lemma 2.7.
[39] Let $r$ be a positive integer, and $f(x)=x^{r}$ be a polynomial over $\mathbb{F}_{q}$.
  1. (1)
    If $\gcd(r,q-1)=1$, then $f(x)$ is a permutation polynomial over $\mathbb{F}_{q}$.
  2. (2)
    If $\gcd(r,q-1)=r$, then $f(x)$ is $r$-to-11 over $\mathbb{F}^{*}_{q}$.
2.3 Arcs in $\mathrm{PG}(3,q)$
An arc in $\mathrm{PG}(3,q)$, $q\geq 4$ has at most $q+1$ points [17, Theorems 21.2.4 and 21.3.8]. If $q=2^{m}$, a $(q+1)$-arc in $\mathrm{PG}(3,q)$ is projectively equivalent to the point-set
$\mathcal{S}_{h}=\{P_{x}=(1,x,x^{2^{h}},x^{2^{h}+1})\mid x\in\mathbb{F}_{q}\}\cup\{P_{\infty}=(0,0,0,1)\},$
where $\gcd(m,h)=1$ [17, Theorem 21.3.15]. For $h=1$, $\mathcal{S}_{1}$ is called a twisted cubic. When $q$ is odd, every $(q+1)$-arc in $\mathrm{PG}(3,q)$ is a twisted cubic. A line in $\mathrm{PG}(3,q)$ joining two distinct points of $\mathcal{S}_{h}$ is called a real chord. Note that no four points of $\mathcal{S}_{h}$ are coplanar. Therefore, each real chord contains exactly two points of $\mathcal{S}_{h}$. Let $R$ denote the set of points lying on real chords but not in $\mathcal{S}_{h}$. Since there are $q(q+1)/2$ real chords, it follows that $|R|=(q^{3}-q)/2$. The planes $\pi(P_{x})$ and $\pi(P_{\infty})$, defined by the equations $x^{2^{h}+1}X_{1}+x^{2^{h}}X_{2}+xX_{3}+X_{4}=0$ and $X_{1}=0$ are called the osculating planes of $\mathcal{S}_{h}$ at $P_{x}$ and $P_{\infty}$, respectively. The osculating plane $\pi(P)$ intersects $\mathcal{S}_{h}$ only at point $P$. We end this section with the following lemma.
Lemma 2.8.
[37, Theorem A] Let $K$ be a $k$-arc of $\mathrm{PG}(3,q)$, $q$ even and $q\neq 2$. Assume that $k>q-\sqrt{q}/2+9/4$. Then $K$ lies on a unique $(q+1)$-arc.
3 NMDS codes of dimension 33
3.1 NMDS codes from hyperovals
When $k=3$, a point-set $\mathcal{S}$ of size $n$ in $\mathrm{PG}(2,q)$ satisfying the properties (N2) and (N3) in Section 1, i.e., $|\mathcal{H}_{u}\cap\mathcal{S}|\leq 3$ for any nonzero vector $u\in\mathbb{F}^{3}_{q}$ and $|\mathcal{H}_{v}\cap\mathcal{S}|=3$ for some nonzero vector $v\in\mathbb{F}^{3}_{q}$, is called an $(n,3)$-arc. Every [n,3,n3][n,3,n-3] NMDS code is equivalent to an $(n,3)$-arc in $\mathrm{PG}(2,q)$. In this section, let $q=2^{m}$ and $\mathbb{F}^{*}_{q}=\mathbb{F}_{q}\setminus\{0\}=\{\alpha_{1},\dots,\alpha_{q-1}\}$, where $m$ is a positive integer. Define $f(x)\in\mathbb{F}_{q}[x]$ to be an o-polynomial, and let $G$ be a $3\times(q+3)$ matrix defined as
$G=(\mathbf{g}_{1},\ldots,\mathbf{g}_{q+2},\mathbf{g}_{q+3})=\begin{pmatrix}1&1&\cdots&1&1&0&0&a\\ \alpha_{1}&\alpha_{2}&\cdots&\alpha_{q-1}&0&0&1&b\\ f(\alpha_{1})&f(\alpha_{2})&\cdots&f(\alpha_{q-1})&0&1&0&c\end{pmatrix}$, (3.1)
where $\mathbf{g}_{q+3}=(a,b,c)^{\mathrm{T}}\neq\mathbf{0}$, and $\mathbf{g}_{q+3}\neq\lambda\mathbf{g}_{i}$ for some $\lambda\in\mathbb{F}^{*}_{q}$ and any $i\in\{1,\dots,q+2\}$. For convenience, we identify each vector $\mathbf{g}_{i}$ with the point in $\mathrm{PG}(2,q)$ spanned by $\mathbf{g}_{i}$. This ensures that the points $\mathbf{g}_{i}$ are distinct for $i=1,\dots,q+3$. Let $\mathcal{C}$ be the linear code over $\mathbb{F}_{q}$ with generator matrix $G$. For any $\mathbf{g}_{q+3}$ , we first prove the following theorem, which generalizes [40, Theorem 8] and [21, Theorem 10].
Theorem 3.9.
Let $q=2^{m}$ with $m\geq 2$. Then $\mathcal{C}$ is a $[q+3,3,q]$ NMDS code over $\mathbb{F}_{q}$ with weight enumerator
$A(z)=1+\frac{(q-1)(q+2)}{2}z^{q}+\frac{q(q-1)(q+2)}{2}z^{q+1}+\frac{q(q-1)}{2}z^{q+2}+\frac{q(q-2)(q-1)}{2}z^{q+3}.$
Proof.
Let $\mathcal{S}_{G}$ denote the set of columns of $G$. Since the first $q+2$ columns of $\mathcal{S}_{G}$ form a hyperoval $\mathcal{HO}$, the condition $|\mathcal{H}_{u}\cap\mathcal{S}_{G}|\leq 3$ holds for any nonzero vector $u\in\mathbb{F}^{3}_{q}$. To show that $\mathcal{S}_{G}$ is an $(n,3)$-arc and thus $\mathcal{C}$ is an NMDS code, we calculate $A_{q}$, i.e., the number of vectors $u=(u_{1},u_{2},u_{3})\in\mathbb{F}^{3}_{q}$ satisfying $|\mathcal{H}_{u}\cap\mathcal{S}_{G}|=3$. Note that if $|\mathcal{H}_{u}\cap\mathcal{S}_{G}|=3$, then we must have $\mathbf{g}_{q+3}\in\mathcal{H}_{u}$. We now consider each possible case for $\mathbf{g}_{q+3}$ in turn.
Case 1. $\mathbf{g}_{q+3}=(a,b,c)^{\mathrm{T}}=(0,1,c)^{\mathrm{T}}$ with $c\in\mathbb{F}^{*}_{q}$.
Then $u_{2}+cu_{3}=0$. If $u_{2}=0$, then $u_{3}=0$ and thus $u=(u_{1},0,0)$. In this case, we have $\mathcal{H}_{u}\cap\mathcal{S}_{G}=\{\mathbf{g}_{q+1},\mathbf{g}_{q+2},\mathbf{g}_{q+3}\}$. If $u_{2}\neq 0$, then $\mathbf{g}_{q+1}$ and $\mathbf{g}_{q+2}$ are not in $\mathcal{H}_{u}$. Suppose that $(1,x,f(x))^{\mathrm{T}}\in\mathcal{H}_{u}$ for some $x\in\mathbb{F}_{q}$. Then $u_{1}+cu_{3}x+u_{3}f(x)=0$, i.e., $f(x)+cx+u_{1}u_{3}^{-1}=0$. By Lemma 2.5, this equation has exactly two solutions and there are q(q1)/2q(q-1)/2 choices for $(u_{1},u_{3})$ satisfying $u_{1}u_{3}^{-1}\in\{f(x)+cx:x\in\mathbb{F}_{q}\}$. Thus, in this case, $A_{q}=(q-1)+q(q-1)/2=(q-1)(q+2)/2$.
Case 2. $\mathbf{g}_{q+3}=(a,b,c)^{\mathrm{T}}=(1,0,c)^{\mathrm{T}}$ with $c\in\mathbb{F}^{*}_{q}$.
Then $u_{1}+cu_{3}=0$. If $u_{1}=0$, then $u_{3}=0$ and thus $u=(0,u_{2},0)$. In this case, we have $\mathcal{H}_{u}\cap\mathcal{S}_{G}=\{\mathbf{g}_{q},\mathbf{g}_{q+1},\mathbf{g}_{q+3}\}$. If $u_{1}\neq 0$, then $\mathbf{g}_{q}$ and $\mathbf{g}_{q+1}$ are not in $\mathcal{H}_{u}$. Moreover, if $u_{2}=0$, then $\mathcal{H}_{u}\cap\mathcal{S}_{G}=\{\mathbf{g}_{j},\mathbf{g}_{q+2},\mathbf{g}_{q+3}\}$, where $f(\alpha_{j})=c$. Finally, we consider the case where $u_{1},u_{2}\neq 0$. Suppose that $(1,x,f(x))^{\mathrm{T}}\in\mathcal{H}_{u}$ for some $x\in\mathbb{F}^{*}_{q}$. Then $cu_{3}+u_{2}x+u_{3}f(x)=0$. By Lemma 2.6, this equation has exactly two solutions and there are (q1)(q2)/2(q-1)(q-2)/2 choices for $(u_{1},u_{2})$. Thus, in this case, $A_{q}=2(q-1)+(q-1)(q-2)/2=(q-1)(q+2)/2$.
Case 3. $\mathbf{g}_{q+3}=(a,b,c)^{\mathrm{T}}=(1,b,0)^{\mathrm{T}}$ with $b\in\mathbb{F}^{*}_{q}$.
This case is similar to the Case 2 and we omit the details.
Case 4. $\mathbf{g}_{q+3}=(a,b,c)^{\mathrm{T}}=(1,b,c)^{\mathrm{T}}$ with $b,c\in\mathbb{F}^{*}_{q}$ and $c\neq f(b)$.
Then $u_{1}+bu_{2}+cu_{3}=0$. If $u_{2}=0$ and $u_{3}\neq 0$ , then $\mathcal{H}_{u}\cap\mathcal{S}_{G}=\{\mathbf{g}_{i},\mathbf{g}_{q+2},\mathbf{g}_{q+3}\}$, where $f(\alpha_{i})=c$. If $u_{3}=0$ and $u_{2}\neq 0$ , then $\mathcal{H}_{u}\cap\mathcal{S}_{G}=\{\mathbf{g}_{j},\mathbf{g}_{q+2},\mathbf{g}_{q+3}\}$, where $\alpha_{j}=b$. Finally, if $u_{2}\neq 0$ and $u_{3}\neq 0$, and $(1,x,f(x))^{\mathrm{T}}\in\mathcal{H}_{u}$ for some $x\in\mathbb{F}^{*}_{q}$, then $u_{1}+u_{2}x+u_{3}f(x)=0$, which can be rewritten as $u_{3}f(x)+u_{2}x+u_{3}c+u_{2}b=0$. By Lemma 2.6, this equation has exactly two solutions and there are (q1)(q2)/2(q-1)(q-2)/2 choices for $(u_{1},u_{2})$. Thus, in this case, $A_{q}=2(q-1)+(q-1)(q-2)/2=(q-1)(q+2)/2$.
From Case 1-Case 4, we see that the number of codewords of weight $q$ in $\mathcal{C}$ is (q1)(q+2)/2(q-1)(q+2)/2 for any possible $\mathbf{g}_{q+3}$. The desired weight enumerator of $\mathcal{C}$ then follows from Lemma 2.2. This completes the proof. ∎
A collineation $\theta(A,\sigma)$ of $\mathrm{PG}(2,q)$ is a projective semilinear map defined by: $(x,y,z)^{\mathrm{T}}\mapsto A(x^{\sigma},y^{\sigma},z^{\sigma})^{\mathrm{T}}$, where $A$ is a nonsingular $3\times 3$ matrix over $\mathbb{F}_{q}$ and $\sigma\in\textup{Aut}(\mathbb{F}_{q})$ (for convenience, we represent points of $\mathrm{PG}(2,q)$ as column vectors). The full collineation group of $\mathrm{PG}(2,q)$ is $\mathrm{P}\Gamma\mathrm{L}(3,q)$. If $\sigma=1$, the collineation is called a homography. All homographies form the group $\mathrm{PGL}(3,q)$. The stabilizer of a hyperoval $\mathcal{HO}$ in $\mathrm{P}\Gamma\mathrm{L}(3,q)$ is called the collineation stabilizer of $\mathcal{HO}$, while its stabilizer in $\mathrm{PGL}(3,q)$ is called the homography stabilizer.
Proposition 3.10.
Let $q=2^{m}$ with $m\geq 2$. Suppose that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are [n,3,n3][n,3,n-3] NMDS codes with generator matrix $G_{1}$ and $G_{2}$ respectively, where $q+2<n<5(q+2)/4$. If
$G_{1}=(\mathbf{g}_{1},\ldots,\mathbf{g}_{q+2},\mathbf{g}_{q+3},\dots,\mathbf{g}_{n})\;\text{and}\;G_{2}=(\mathbf{g}^{\prime}_{1},\ldots,\mathbf{g}^{\prime}_{q+2},\mathbf{g}^{\prime}_{q+3},\dots,\mathbf{g}^{\prime}_{n}),$
where the point-sets $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{q+2}\}$ and $\{\mathbf{g}^{\prime}_{1},\ldots,\mathbf{g}^{\prime}_{q+2}\}$ are hyperovals $\mathcal{HO}_{1}$ and $\mathcal{HO}_{2}$ in $\mathrm{PG}(2,q)$ respectively, then $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are monomially equivalent if and only if there exists a homography $\theta$ such that $\mathcal{HO}^{\theta}_{1}=\mathcal{HO}_{2}$ and the sets of points $\{\mathbf{g}_{q+3}^{\theta},\dots,\mathbf{g}_{n}^{\theta}\}$ and $\{\mathbf{g}^{\prime}_{q+3},\dots,\mathbf{g}^{\prime}_{n}\}$ are equal.
Proof.
Recall that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are monomially equivalent if and only if there exist a matrix $L\in\mathrm{GL}(k,q)$ and an $n\times n$ monomial matrix $M$ such that $LG_{1}M=G_{2}$. This is equivalent to the sets of projective points
$\{L\mathbf{g}_{1},\ldots,L\mathbf{g}_{q+2},L\mathbf{g}_{q+3},\dots,L\mathbf{g}_{n}\}\;\text{and}\;\{\mathbf{g}^{\prime}_{1},\ldots,\mathbf{g}^{\prime}_{q+2},\mathbf{g}^{\prime}_{q+3},\dots,\mathbf{g}^{\prime}_{n}\}$
are equal. We can write $L\mathbf{g}_{i}=\mathbf{g}_{i}^{\theta}$ for some homography $\theta$, where $i=1,\dots,q+2$. Since $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{q+2}\}$ is a hyperoval $\mathcal{HO}$ in $\mathrm{PG}(2,q)$, the set $\mathcal{HO}^{\theta}_{1}:=\{\mathbf{g}^{\theta}_{1},\ldots,\mathbf{g}^{\theta}_{q+2}\}$ is also a hyperoval. Given $n<5(q+2)/4$, it follows that
$|\mathcal{HO}_{2}\cap\mathcal{HO}^{\theta}_{1}|\geq q+2-2(n-q-2)>(q+2)/2.$
By Lemma 2.4, we conclude that $\mathcal{HO}_{2}=\mathcal{HO}^{\theta}_{1}$, and thus $\{\mathbf{g}_{q+3}^{\theta},\dots,\mathbf{g}_{n}^{\theta}\}=\{\mathbf{g}^{\prime}_{q+3},\dots,\mathbf{g}^{\prime}_{n}\}$. This completes the proof. ∎
Corollary 3.11.
Let $\mathscr{C}$ be the set of [n,3,n3][n,3,n-3] NMDS codes over $\mathbb{F}_{q}$ with $q+2<n<5(q+2)/4$, such that the first $q+2$ columns of their generator matrices form a hyperoval $\mathcal{HO}$ in $\mathrm{PG}(2,q)$. Then the number of monomial equivalence classes of NMDS codes in $\mathscr{C}$ is equal to the number of orbits of the homography stabilizer of $\mathcal{HO}$ acting on the set formed by the remaining nq2n-q-2 columns of the generator matrices.
3.2 Monomial equivalence classes of NMDS codes
Recall that a $[q+3,3,q]$ NMDS code was constructed in Theorem 3.9. In this subsection, for the cases where $f(x)$ in (3.1) corresponds to different o-polynomials, we determine the number of monomial equivalence classes for the corresponding NMDS codes generated by (3.1). Let $\mathcal{HO}$ denote the hyperoval corresponding to $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{q+2}\}$ in (3.1) and $M$ be the homography stabilizer of $\mathcal{HO}$.
Lemma 3.12.
[16, Corollary 6][28, Theorem 4.11.4] A regular hyperoval in $\mathrm{PG}(2,q)$, $q>4$ even, has homography stabilizer of order $q(q^{2}-1)$ isomorphic to $\mathrm{PGL}(2,q)$. This stabilizer consists of the maps: $(x,y,z)^{\mathrm{T}}\mapsto A(x,y,z)^{\mathrm{T}}$, where
$\left.A=\left(\begin{array}[]{ccc}a&0&c\\ \sqrt{ab}&1&\sqrt{cd}\\ b&0&d\end{array}\right.\right),\mathrm{~and~}ad+bc=1.$
Theorem 3.13.
Let $\mathscr{C}$ be the set of NMDS codes over $\mathbb{F}_{q}$ with generator matrices of the form (3.1). If $f(x)$ is the regular o-polynomial, i.e., $f(x)=x^{2}$, then there is only one equivalence class in $\mathscr{C}$.
Proof.
We provide details only for $q\geq 8$, as the cases $q=2$ or $4$ can be verified using Magma [7]. Note that the generator matrices of all NMDS codes in $\mathscr{C}$ differ only in the last column vector $\mathbf{g}_{q+3}$. To determine the number of equivalence classes in $\mathscr{C}$, it suffices to consider the $q^{2}-1$ cases where $\mathbf{g}_{q+3}$ is different from the points in $\mathcal{HO}$. Let $\mathcal{C}$ be the NMDS code in $\mathscr{C}$ with $\mathbf{g}_{q+3}=(0,1,1)^{\mathrm{T}}$. Recall that $M$ is the homography stabilizer of $\mathcal{HO}$. We regard $\mathbf{g}_{q+3}$ as a point in $\mathrm{PG}(2,q)$ and compute the stabilizer $M_{\mathbf{g}_{q+3}}$ of $\mathbf{g}_{q+3}$ in $M$. By Lemma 3.12, for any $\theta\in M$, we have
$\mathbf{g}_{q+3}^{\theta}=\left(\begin{array}[]{ccc}a&0&c\\ \sqrt{ab}&1&\sqrt{cd}\\ b&0&d\end{array}\right)\begin{pmatrix}0\\ 1\\ 1\end{pmatrix}=\begin{pmatrix}c\\ 1+\sqrt{cd}\\ d\end{pmatrix}$
for some $a,b,c,d\in\mathbb{F}_{q}$ satisfying $ad+bc=1$. Suppose that $\theta$ stabilizes $\mathbf{g}_{q+3}$. Then $c=0,d=1$. Since $ad+bc=1$, we have $a=1$. It follows from $b\in\mathbb{F}_{q}$ that $|M_{\mathbf{g}_{q+3}}|=q$, and thus $|\mathbf{g}_{q+3}^{M}|=q^{2}-1$. By Corollary 3.11, we conclude that all NMDS codes in $\mathscr{C}$ are monomially equivalent. This completes the proof. ∎
Remark 3.14.
For regular o-polynomial, by setting $\mathbf{g}_{q+3}$ as $(0,1,1)^{\mathrm{T}}$ and $(1,1,0)^{\mathrm{T}}$ in (3.1), we obtain two distinct matrices $G_{1}$ and $G_{2}$. Two $[q+3,3,q]$ NMDS codes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ with generator matrices $G_{1}$ and $G_{2}$, were constructed in [40, Theorem 8] and [21, Theorem 10], respectively. By Theorem 3.13, we see that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are monomially equivalent.
Two [q+1,3,q2][q+1,3,q-2] NMDS codes $\mathcal{C}_{3}$ and $\mathcal{C}_{4}$ with the same weight enumerator were provided in [40, Theorem 10] and [21, Theorem 13] respectively, where $q=2^{m}$ and $m\geq 3$ is odd. The generator matrices $G_{i}$ of $\mathcal{C}_{i}$, for $i=3,4$, are as follows:
$G_{3}=\begin{pmatrix}1&1&\cdots&1&1&1\\ \alpha_{1}&\alpha_{2}&\cdots&\alpha_{q-1}&1&0\\ \alpha^{2}_{1}&\alpha^{2}_{2}&\cdots&\alpha^{2}_{q-1}&0&1\end{pmatrix},\;G_{4}=\begin{pmatrix}1&1&\cdots&1&1&0\\ \alpha_{1}&\alpha_{2}&\cdots&\alpha_{q-1}&1&1\\ \alpha^{2}_{1}&\alpha^{2}_{2}&\cdots&\alpha^{2}_{q-1}&0&1\end{pmatrix}$,
where $\mathbb{F}^{*}_{q}=\mathbb{F}_{q}\setminus\{0\}=\{\alpha_{1},\dots,\alpha_{q-1}\}$.
Theorem 3.15.
Linear codes $\mathcal{C}_{3}$ and $\mathcal{C}_{4}$ are not monomially equivalent.
Proof.
Note that the first $q$ columns of $G_{3}$ and $G_{4}$ are identical and the first q1q-1 columns form a (q1)(q-1)-arc $\mathcal{A}=\mathcal{HO}\setminus\mathcal{S}$, where $\mathcal{HO}$ is a regular hyperoval and $\mathcal{S}=\{(1,0,0)^{\mathrm{T}},(0,0,1)^{\mathrm{T}},(0,1,0)^{\mathrm{T}}\}$. Let $X$ and $Y$ denote the sets of points $\{(1,1,0)^{\mathrm{T}},(1,0,1)^{\mathrm{T}}\}$ and $\{(1,1,0)^{\mathrm{T}},(0,1,1)^{\mathrm{T}}\}$, respectively. Suppose that $\mathcal{C}_{3}$ and $\mathcal{C}_{4}$ are monomially equivalent. Similarly to the proof of Proposition 3.10, there exists a homography $\theta$ such that $\mathcal{A}^{\theta}\cup X^{\theta}=\mathcal{A}\cup Y$. Note that
$|\mathcal{HO}\cap\mathcal{HO}^{\theta}|=|(\mathcal{A}\cup\mathcal{S})\cap(\mathcal{A}^{\theta}\cup\mathcal{S}^{\theta})|\geq|\mathcal{A}\cap\mathcal{A}^{\theta}|\geq q-5>(q+2)/2$
for $q\geq 32$. By Lemma 2.4, this implies that $\mathcal{HO}=\mathcal{HO}^{\theta}$, and consequently $\mathcal{A}=\mathcal{A}^{\theta},\mathcal{S}=\mathcal{S}^{\theta}$. By Lemma 3.12, the matrix representing $\theta$ is of the form $\left(\begin{array}[]{ccc}a&0&c\\ \sqrt{ab}&1&\sqrt{cd}\\ b&0&d\end{array}\right)$, where $ad+bc=1$. Since $\mathcal{S}^{\theta}=\mathcal{S}$, we have
$\{(1,0,0)^{\mathrm{T}},(0,0,1)^{\mathrm{T}},(0,1,0)^{\mathrm{T}}\}=\{(a,\sqrt{ab},b)^{\mathrm{T}},(c,\sqrt{cd},d)^{\mathrm{T}},(0,1,0)^{\mathrm{T}}\}.$
It follows that either $a=d=1,b=c=0$ or $a=d=0,b=c=1$. In either case, $\theta$ maps $(1,0,1)^{\mathrm{T}}$ to $(a+c,\sqrt{ab}+\sqrt{cd},b+d)=(1,0,1)^{\mathrm{T}}\notin Y$, which is a contradiction.
Next, we consider the case $q=8$. Note that if $\mathcal{C}_{3}$ and $\mathcal{C}_{4}$ are equivalent, then their extended codes generated by matrices obtained by adding the same columns must also be equivalent. Let $\omega$ be a primitive element of $\mathbb{F}^{*}_{8}$. We added the column $(1,\omega,1)^{\mathrm{T}}$ to $G_{3}$ and $G_{4}$ to generate two linear codes $\mathcal{C}^{\prime}_{3}$ and $\mathcal{C}^{\prime}_{4}$, respectively. Using Magma [7], we verify that $\mathcal{C}^{\prime}_{3}$ is a $[10,3,7]$ linear code with weight enumerator
$A(z)=1+63z^{7}+126z^{8}+189z^{9}+133z^{10}$
and $\mathcal{C}^{\prime}_{4}$ is a $[10,3,7]$ linear code with weight enumerator
$A(z)=1+56z^{7}+147z^{8}+168z^{9}+140z^{10}.$
Since the weight enumerators of $\mathcal{C}^{\prime}_{3}$ and $\mathcal{C}^{\prime}_{4}$ differ, $\mathcal{C}^{\prime}_{3}$ and $\mathcal{C}^{\prime}_{4}$ are not equivalent. This implies that $\mathcal{C}_{3}$ and $\mathcal{C}_{4}$ are also not equivalent. This completes the proof. ∎
Note that every hyperoval is regular for $q=2,4,8$. From now on, we suppose that $q=2^{m},m\geq 4$.
Lemma 3.16.
[25, Theorem 4.3][28, Theorem 4.11.6] An irregular translation hyperoval with o-polynomial $f(x)=x^{\alpha}$ in $\mathrm{PG}(2,q)$, where $\alpha=2^{h},q=2^{m}\geq 16$, and $\gcd(h,m)=1$, has homography stabilizer of order q(q1)q(q-1). This stabilizer is isomorphic to $\mathrm{AGL}(1,q)$ and consists of the maps: $(x,y,z)^{\mathrm{T}}\mapsto A(x,y,z)^{\mathrm{T}}$, where
$\left.A=\left(\begin{array}[]{ccc}1&0&0\\ a&b&0\\ a^{\alpha}&0&b^{\alpha}\end{array}\right.\right),\mathrm{~with~}a,b\in\mathbb{F}_{q},b\neq 0.$
Theorem 3.17.
Let $\mathscr{C}$ be the set of NMDS codes over $\mathbb{F}_{q}$ with generator matrices of the form (3.1). If $f(x)=x^{\alpha}$ is an irregular translation o-polynomial, where $\alpha=2^{h},q=2^{m}\geq 16$, and $\gcd(h,m)=1$, then there are two equivalence classes in $\mathscr{C}$.
Proof.
Similarly to the proof of Theorem 3.13, let $\mathcal{C}_{x,k}$ be the NMDS code in $\mathscr{C}$ with $\mathbf{g}_{q+3}=(1,x,x^{\alpha}+k)^{\mathrm{T}}$, where $x\in\mathbb{F}_{q}$ and $k\in\mathbb{F}^{*}_{q}$. By Lemma 3.16, for any $\theta\in M$, we have
$\mathbf{g}_{q+3}^{\theta}=\left(\begin{array}[]{ccc}1&0&0\\ a&b&0\\ a^{\alpha}&0&b^{\alpha}\end{array}\right)\begin{pmatrix}1\\ x\\ x^{\alpha}+k\end{pmatrix}=\begin{pmatrix}1\\ a+bx\\ (a+bx)^{\alpha}+kb^{\alpha}\end{pmatrix}$
for some $a\in\mathbb{F}_{q}$ and $b\in\mathbb{F}^{*}_{q}$. Suppose that $\theta$ stabilizes $\mathbf{g}_{q+3}$. Then $a+bx=x$ and $b^{\alpha}=1$. Since $f(x)=x^{\alpha}$ is a permutation polynomial over $\mathbb{F}_{q}$, it follows that $a=0,b=1$. Consequently, $|M_{\mathbf{g}_{q+3}}|=1$ and so $s_{1}:=|\mathbf{g}_{q+3}^{M}|=q^{2}-q$.
Let $\mathcal{C}^{u}$ be the NMDS code in $\mathscr{C}$ with $\mathbf{g}_{q+3}=(0,1,u)^{\mathrm{T}}$, where $u\in\mathbb{F}^{*}_{q}$. By Lemma 3.16, for any $\theta\in M$, we have
$\mathbf{g}_{q+3}^{\theta}=\left(\begin{array}[]{ccc}1&0&0\\ a&b&0\\ a^{\alpha}&0&b^{\alpha}\end{array}\right)\begin{pmatrix}0\\ 1\\ u\end{pmatrix}=\begin{pmatrix}0\\ b\\ ub^{\alpha}\end{pmatrix}$
for some $a\in\mathbb{F}_{q}$ and $b\in\mathbb{F}^{*}_{q}$. Suppose that $\theta$ stabilizes $\mathbf{g}_{q+3}$. Then $b^{2^{h}-1}=b^{\alpha-1}=1$. Since $b^{q-1}=b^{2^{m}-1}=1$, we have $b^{\gcd(2^{h}-1,2^{m}-1)}=1$ and so $b=1$. It follows from $a\in\mathbb{F}_{q}$ that $|M_{\mathbf{g}_{q+3}}|=q$ and so $s_{2}:=|\mathbf{g}_{q+3}^{M}|=q-1$. Since $s_{1}+s_{2}=q^{2}-1$, by Corollary 3.11, we conclude that there are two equivalence classes in $\mathscr{C}$. This completes the proof. ∎
Lemma 3.18.
[25, Theorem 4.4][28, Theorem 4.12.8] Let $\mathcal{HO}$ be a hyperoval with o-polynomial $f(x)=x^{n}$, but not a translation hyperoval. If $n^{2}-n+1\not\equiv 0\pmod{q-1}$, then $\mathcal{HO}$ has homography stabilizer $H$ of order (q1)(q-1) which consists of the maps: $(x,y,z)^{\mathrm{T}}\mapsto A(x,y,z)^{\mathrm{T}}$, where
$\left.A=\left(\begin{array}[]{ccc}1&0&0\\ 0&a&0\\ 0&0&a^{n}\end{array}\right.\right),\mathrm{~with~}a\in\mathbb{F}^{*}_{q}.$
If $n^{2}-n+1\equiv 0\pmod{q-1}$, then $H$ contains an additional homography $\sigma$ such that |H|=3(q1)|H|=3(q-1) and the matrix representing $\sigma$ is of the form $\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)$.
Theorem 3.19.
Let $\mathscr{C}$ be the set of NMDS codes over $\mathbb{F}_{q}$ with generator matrices of the form (3.1). If $f(x)=x^{n}$ is an o-polynomial but not a translation o-polynomial, then there are $(q+1)/3$ equivalence classes in $\mathscr{C}$ if $n^{2}-n+1\equiv 0\pmod{q-1}$. Otherwise, there are $q+r$ equivalence classes in $\mathscr{C}$, where $r:=\#\{a\in\mathbb{F}^{*}_{q}:a^{n-1}=1\}$.
Proof.
We first consider the case where $n^{2}-n+1\not\equiv 0\pmod{q-1}$. Similarly to the proof of Theorem 3.13, let $\mathcal{C}_{x,k}$ be the NMDS code in $\mathscr{C}$ with $\mathbf{g}_{q+3}=(1,x,x^{n}+k)^{\mathrm{T}}$, where $x\in\mathbb{F}_{q}$ and $k\in\mathbb{F}^{*}_{q}$. By Lemma 3.18, for any $\theta_{a}\in M$, we have
$\mathbf{g}_{q+3}^{\theta_{a}}=\left(\begin{array}[]{ccc}1&0&0\\ 0&a&0\\ 0&0&a^{n}\end{array}\right)\begin{pmatrix}1\\ x\\ x^{n}+k\end{pmatrix}=\begin{pmatrix}1\\ ax\\ (ax)^{n}+a^{n}k\end{pmatrix}$
for some $a\in\mathbb{F}^{*}_{q}$. Suppose that $\theta_{a}$ stabilizes $\mathbf{g}_{q+3}$. Then $a=1$. It follows that $|M_{\mathbf{g}_{q+3}}|=1$, and thus $|\mathbf{g}_{q+3}^{M}|=q-1$. Hence, by Corollary 3.11, there are $q$ equivalence classes of NMDS codes in $\mathscr{C}$ such that $\mathbf{g}_{q+3}$ is of the form $(1,x,x^{n}+k)^{\mathrm{T}}$, where $x\in\mathbb{F}_{q}$ and $k\in\mathbb{F}^{*}_{q}$. Let $\mathcal{C}^{u}$ be the NMDS code in $\mathscr{C}$ with $\mathbf{g}_{q+3}=(0,1,u)^{\mathrm{T}}$, where $u\in\mathbb{F}^{*}_{q}$. By Lemma 3.18, for any $\theta_{a}\in M$, we have
$\mathbf{g}_{q+3}^{\theta}=\left(\begin{array}[]{ccc}1&0&0\\ 0&a&0\\ 0&0&a^{n}\end{array}\right)\begin{pmatrix}0\\ 1\\ u\end{pmatrix}=\begin{pmatrix}0\\ a\\ ua^{n}\end{pmatrix}$
for some $a\in\mathbb{F}^{*}_{q}$. Suppose that $\theta_{a}$ stabilizes $\mathbf{g}_{q+3}$. Then $a^{n-1}=1$. Let $r:=\#\{a\in\mathbb{F}^{*}_{q}:a^{n-1}=1\}$. It follows that $|M_{\mathbf{g}_{q+3}}|=r$ and so $|\mathbf{g}_{q+3}^{M}|=(q-1)/r$. By Corollary 3.11, there are $r$ equivalence classes of NMDS codes in $\mathscr{C}$ such that $\mathbf{g}_{q+3}$ is of the form $(0,1,u)^{\mathrm{T}}$, where $u\in\mathbb{F}^{*}_{q}$. Hence, there are $q+r$ equivalence classes in $\mathscr{C}$.
Next, we consider the case where $n^{2}-n+1\equiv 0\pmod{q-1}$. By Lemma 3.18, there exists a homography $\sigma$ represented by the matrix $\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)$. We claim that $\sigma,\sigma\theta_{a}$ and $\sigma^{2}\theta_{a}$ cannot stabilize any point not in $\mathcal{HO}$ for any $a\in\mathbb{F}^{*}_{q}$. We only provide the details for $\sigma\theta_{a}$ and the other cases are similar. It is obvious that the points of the form $(0,1,u)^{\mathrm{T}}$ cannot stabilized by $\sigma\theta_{a}$ for any $a\in\mathbb{F}^{*}_{q}$, where $u\in\mathbb{F}^{*}_{q}$. Suppose that $\mathbf{g}_{q+3}=(1,x,x^{n}+k)^{\mathrm{T}}$ is stabilized by $\sigma\theta_{a}$ for some $a\in\mathbb{F}^{*}_{q}$, where $x$, $k$ and $x^{n}+k$ are all nonzero. Since
$\mathbf{g}_{q+3}^{\sigma\theta_{a}}=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)\left(\begin{array}[]{ccc}1&0&0\\ 0&a&0\\ 0&0&a^{n}\end{array}\right)\begin{pmatrix}1\\ x\\ x^{n}+k\end{pmatrix}=\begin{pmatrix}ax\\ a^{n}(x^{n}+k)\\ 1\end{pmatrix}$,
we have $\frac{ax}{1}=\frac{a^{n}(x^{n}+k)}{x}=\frac{1}{x^{n}+k}$ and so $x^{n}+k=(ax)^{-1}$ and $x^{3}=a^{n-2}$. From $n^{2}-n+1\equiv 0\pmod{q-1}$, it follows that $x^{3(n+1)}=a^{(n-2)(n+1)}=a^{-3}$. Since $\gcd(3,q-1)=1$, we obtain $a=x^{-(n+1)}$. Thus $x^{n}+k=(ax)^{-1}=x^{n}$ and so $k=0$, which is a contradiction. On the other hand, we have $\gcd(n-1,q-1)=1$ by $n^{2}-n+1\equiv 0\pmod{q-1}$. Thus $r:=\#\{a\in\mathbb{F}^{*}_{q}:a^{n-1}=1\}=1$. Combining this with the proof in the first paragraph, we conclude that there are $(q+1)/3$ equivalence classes in $\mathscr{C}$. This completes the proof.∎
Theorem 3.20.
Let $\mathscr{C}$ be the set of NMDS codes over $\mathbb{F}_{q}$ with generator matrices of the form (3.1). If $f(x)$ is the Cherowitzo o-polynomial, then there are $q^{2}-1$ equivalence classes in $\mathscr{C}$.
Proof.
By [5, Corollary 4.4], the Cherowitzo hyperoval has trivial homography stabilizer. The claim then follows from Corollary 3.11. ∎
Theorem 3.21.
Let $\mathscr{C}$ be the set of NMDS codes over $\mathbb{F}_{q}$ with generator matrices of the form (3.1). If $f(x)$ is a Payne o-polynomial or Adelaide o-polynomial, then there are $(q^{2}+q-2)/2$ equivalence classes in $\mathscr{C}$.
Proof.
By [38], the Payne hyperoval has homography stabilizer of order $2$, generated by the following matrix $\left(\begin{array}[]{ccc}0&1&0\\ 1&0&0\\ 0&0&1\end{array}\right)$. This stabilizer fixes points of the form $(1,1,m)^{\mathrm{T}}$, where $m\in\mathbb{F}_{q}\setminus\{1\}$, and exchanges the remaining $q^{2}-q$ points not in the hyperoval pairwise. The claim then follows from Corollary 3.11.
By [31], the Adelaide hyperoval has homography stabilizer of order $2$, generated by the following matrix $\left(\begin{array}[]{ccc}1&0&0\\ \beta+\beta^{-1}&1&0\\ 0&0&1\end{array}\right)$, where $\beta\in\mathbb{F}_{q^{2}}\setminus\{1\}$ satisfies $\beta^{q+1}=1$. This stabilizer fixes points of the form $(0,1,m)^{\mathrm{T}}$, where $m\in\mathbb{F}^{*}_{q}$ and exchanges the remaining $q^{2}-q$ points not in the hyperoval pairwise. The claim then follows from Corollary 3.11. ∎
For small $q$, each Subiaco hyperoval falls into one of the previously known classes of hyperovals. If $q=16$, a Subiaco hyperoval is a Lunelli-Sce hyperoval [22]. Its homography stabilizer has order $36$ and is isomorphic to $\mathrm{C}^{3}_{2}\rtimes\mathrm{C}_{4}$ [29]. If $q=32$, a Subiaco hyperoval is a Payne hyperoval [27]. If $q=64$, the two projectively distinct Subiaco hyperovals are the two Penttila-Pinneri irregular hyperovals [32]. The homography stabilizer is either $\mathrm{C}_{5}$ or $\mathrm{D}_{10}$. If $q=2^{m}>64$ with $m\not\equiv 2\pmod{4}$, the homography stabilizer of a Subiaco hyperoval is a cyclic group of order $2$, similar to the case $q=32$ [26]. If $q=2^{m}>64$ with $m\equiv 2\pmod{4}$, the homography stabilizer of a Subiaco hyperoval is either $\mathrm{C}_{5}$ or $\mathrm{D}_{10}$, similar to the case $q=64$ [30]. We provide details for only one case.
Theorem 3.22.
Let $\mathscr{C}$ be the set of NMDS codes over $\mathbb{F}_{q}$ with generator matrices of the form (3.1), where $q=2^{m}>64$ with $m\equiv 2\pmod{4}$. If $f(x)=\frac{\delta^{2}(x^{4}+x)}{\left(x^{2}+\delta x+1\right)^{2}}+x^{1/2}$ is a Subiaco o-polynomial with $\delta^{2}+\delta+1=0$, then there are $(q^{2}-1)/5$ equivalence classes in $\mathscr{C}$.
Proof.
In this case, the Subiaco hyperoval has homography stabilizer $\mathrm{C}_{5}$, generated by $\theta:\begin{aligned} (x,y,z)^{\mathrm{T}}\mapsto(y,x+\delta y,z+\delta^{2}y)^{\mathrm{T}}\end{aligned}$. We can verify that the points not in the Subiaco hyperoval are not stabilized by $\theta$. The claim then follows from Corollary 3.11. ∎
Theorem 3.23.
Let $\mathscr{C}$ be the set of NMDS codes over $\mathbb{F}_{q}$ with generator matrices of the form (3.1). If $f(x)$ is the O’Keefe-Penttila o-polynomial, then there are $341$ equivalence classes in $\mathscr{C}$.
Proof.
By [24, Theroem 3.2], the O’Keefe-Penttila hyperoval has homography stabilizer of order $3$, generated by the matrix $\left(\begin{array}[]{ccc}0&1&0\\ 1&1&1\\ 0&0&1\end{array}\right)$. It cannot fix any point not in the hyperoval. The claim then follows from Corollary 3.11. ∎
4 NMDS codes of dimension 44
Recall that the point-set
$\mathcal{S}_{h}=\{P_{x}=(1,x,x^{2^{h}},x^{2^{h}+1})\mid x\in\mathbb{F}_{q}\}\cup\{P_{\infty}=(0,0,0,1)\}$
is a $(q+1)$-arc in $\mathrm{PG}(3,q)$ for some prime power $q=p^{m}$, where $\gcd(m,h)=1$. Let the columns of matrix $G$ be constructed by the $q+1$ points of $\mathcal{S}_{h}$ and one point not in $\mathcal{S}_{h}$. Then $G$ has the form:
$G=(\mathbf{g}_{1},\ldots,\mathbf{g}_{q+1},\mathbf{g}_{q+2})=\begin{pmatrix}1&1&\cdots&1&1&0&a\\ \alpha_{1}&\alpha_{2}&\cdots&\alpha_{q-1}&0&0&b\\ \alpha_{1}^{2^{h}}&\alpha_{2}^{2^{h}}&\cdots&\alpha_{q-1}^{2^{h}}&0&0&c\\ \alpha_{1}^{2^{h}+1}&\alpha_{2}^{2^{h}+1}&\cdots&\alpha_{q-1}^{2^{h}+1}&0&1&d\end{pmatrix}$, (4.1)
where $\mathbb{F}^{*}_{q}=\mathbb{F}_{q}\setminus\{0\}=\{\alpha_{1},\dots,\alpha_{q-1}\}$ and $\mathbf{g}_{q+2}=(a,b,c,d)^{\mathrm{T}}\neq\mathbf{0}$. Let $\mathcal{S}_{G}$ denote the set of points $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{q+1},\mathbf{g}_{q+2}\}$.
4.1 $q$ even
In this subsection, we suppose that $q=2^{m}\geq 8$ such that $\gcd(h,m)=1$. Let $K_{h}$ be the stabilizer of $\mathcal{S}_{h}$ in $\mathrm{PGL}(4,q)$. Then $K_{h}$ is isomorphic to $\mathrm{PGL}(2,q)$, and its elements are represented by the matrices
$M_{a,b,c,d}=\left(\begin{array}[]{cccc}a^{2^{h}+1}&a^{2^{h}}c&ac^{2^{h}}&c^{2^{h}+1}\\ a^{2^{h}}b&a^{2^{h}}d&bc^{2^{h}}&c^{2^{h}}d\\ ab^{2^{h}}&b^{2^{h}}c&ad^{2^{h}}&cd^{2^{h}}\\ b^{2^{h}+1}&b^{2^{h}}d&bd^{2^{h}}&d^{2^{h}+1}\end{array}\right)\text{, }$
where $a,b,c,d\in\mathbb{F}_{q}$ satisfy $ad+bc=1$, see [17, Lemma 21.3.13, Theorem 21.3.17]. We now present the results on the orbits of points and planes under the action of $K_{h}$.
Lemma 4.24.
[8, Proposition 3.2] The group $K_{h}$ has five orbits on the points of $\mathrm{PG}(3,q)$:
  1. (1)
    $\mathcal{M}_{1}:=\mathcal{S}_{h}$ of size $q+1$;
  2. (2)
    $\mathcal{M}_{2}$ of size $q^{2}+q$, consisting of points on exactly two osculating planes;
  3. (3)
    $\mathcal{M}_{3}$ of size $(q^{3}-q)/6$, consisting of points on exactly three osculating planes;
  4. (4)
    $\mathcal{M}_{4}$ of size $(q^{3}-q)/2$, consisting of points on exactly one osculating planes;
  5. (5)
    $\mathcal{M}_{5}$ of size $(q^{3}-q)/3$, consisting of points not contained in any osculating planes.
Let $R$ be the set of points on real chords but not in $\mathcal{S}_{h}$. If $q\equiv-1\pmod{3}$, then $R=\mathcal{M}_{4}$; if $q\equiv 1\pmod{3}$, then $R=\mathcal{M}_{3}\cup\mathcal{M}_{5}$.
Recall that the osculating planes of $\mathcal{S}_{h}$ at $P_{x}$ and $P_{\infty}$ are defined by the equations $x^{2^{h}+1}X_{1}+x^{2^{h}}X_{2}+xX_{3}+X_{4}=0$ and $X_{1}=0$, respectively. Let $\pi(P)$ denote the osculating planes of $\mathcal{S}_{h}$ at $P$ and $\pi(\mathcal{M}_{i})=\{\pi(P)|P\in\mathcal{M}_{i}\}$, where $i=1,\dots,5$. The $K_{h}$-orbits on the planes of $\mathrm{PG}(3,q)$ can be described by the following lemma.
Lemma 4.25.
[8, Proposition 3.4] The group $K_{h}$ has five orbits on the planes of $\mathrm{PG}(3,q)$:
  1. (1)
    $\pi(\mathcal{M}_{1})$ consisting of the $q+1$ osculating planes;
  2. (2)
    $\pi(\mathcal{M}_{2})$ of size $q^{2}+q$, consisting of planes containing exactly two points of $\mathcal{S}_{h}$;
  3. (3)
    $\pi(\mathcal{M}_{3})$ of size $(q^{3}-q)/6$, consisting of planes containing exactly three points of $\mathcal{S}_{h}$;
  4. (4)
    $\pi(\mathcal{M}_{4})$ of size $(q^{3}-q)/2$, consisting of planes containing exactly one point of $\mathcal{S}_{h}$;
  5. (5)
    $\pi(\mathcal{M}_{5})$ of size $(q^{3}-q)/3$, consisting of planes containing no point of $\mathcal{S}_{h}$.
Recall that a [q+2,4,q2][q+2,4,q-2] NMDS code can be viewed as a point-set $\mathcal{S}$ of size $q+2$ in $\mathrm{PG}(3,q)$ satisfying the following conditions:
  1. (1)
    any three points of $\mathcal{S}$ generate a plane in $\mathrm{PG}(3,q)$;
  2. (2)
    there exist four points in $\mathcal{S}$ that lie on a plane in $\mathrm{PG}(3,q)$;
  3. (3)
    every five points of $\mathcal{S}$ generate $\mathrm{PG}(3,q)$.
Let $\mathcal{C}$ be the linear code over $\mathbb{F}_{q}$ with the generator matrix $G$ defined in (4.1). Suppose that the point $\mathbf{g}_{q+2}=(a,b,c,d)^{\mathrm{T}}$ does not lie on any real chord. Then any three columns of matrix $G$ are linearly independent, ensuring that condition (1) holds. Since the first $q+1$ columns of $G$ constitute the $(q+1)$-arc $\mathcal{S}_{h}$, we have $|\mathcal{H}_{u}\cap\mathcal{S}_{h}|\leq 3$ for any plane $\mathcal{H}_{u}$. It follows that $|\mathcal{H}_{u}\cap\mathcal{S}_{G}|\leq 4$, which implies that condition (3) holds. We now verify condition (2) for different cases and provide some equivalence classes of NMDS codes.
Theorem 4.26.
Let $\mathscr{C}$ be the set of linear codes over $\mathbb{F}_{q}$ with generator matrices of the form (4.1). Then there are either two or three equivalence classes of NMDS codes in $\mathscr{C}$ depending on whether $q\equiv 1\pmod{3}$ or $q\equiv-1\pmod{3}$. Moreover, the weight distributions of these codes are provided in Table 2.
Proof.
By Lemma 4.24, if $q\equiv-1\pmod{3}$ and $\mathbf{g}_{q+2}\in\mathcal{M}_{4}$ or $q\equiv 1\pmod{3}$ and $\mathbf{g}_{q+2}\in\mathcal{M}_{3}\cup\mathcal{M}_{5}$, then the corresponding linear code is not an NMDS code. For the remaining cases, we only need consider the number of nonzero vectors $u\in\mathbb{F}^{4}_{q}$ such that $|\mathcal{H}_{u}\cap\mathcal{S}_{G}|=4$. Suppose that a nonzero vector $u\in\mathbb{F}^{4}_{q}$ satisfies this condition. Then $\mathbf{g}_{q+2}\in\mathcal{H}_{u}$ and so $|\mathcal{H}_{u}\cap\mathcal{S}_{h}|=3$. By Lemma 4.25, all planes containing exactly three points of $\mathcal{S}_{h}$ form a $K_{h}$-orbit $\pi(\mathcal{M}_{3})$. This means the number of nonzero vectors $u\in\mathbb{F}^{4}_{q}$ such that $|\mathcal{H}_{u}\cap\mathcal{S}_{G}|=4$ is equal to q1q-1 times the number of planes in $\pi(\mathcal{M}_{3})$ through the point $\mathbf{g}_{q+2}$. By [8, Table 1], these numbers are all great than 0 and only depend on the $K_{h}$-orbit on points to which $\mathbf{g}_{q+2}$ belongs. Thus, the remaining linear codes are all NMDS codes. Obviously, the codes corresponding to $\mathbf{g}_{q+2}$ in the same $K_{h}$-orbit are monomially equivalent, while the codes corresponding to $\mathbf{g}_{q+2}$ in different $K_{h}$-orbits are not monomially equivalent as they have different numbers of minimum weight codes. Therefore, there are two or three equivalence classes of NMDS codes in $\mathscr{C}$ depending on whether $q\equiv 1\pmod{3}$ or $q\equiv-1\pmod{3}$. Finally, by Lemma 2.2, the weight distributions of these codes can be fully determined and we provide them in Table 2. ∎
Table 2: The weight distributions of NMDS codes in Theorem 4.26 when $\mathbf{g}_{q+2}\in\mathcal{M}_{i}$.
$\mathcal{M}_{2}$ $\mathcal{M}_{3}$ $\mathcal{M}_{4}$ $\mathcal{M}_{5}$
$A_{q-2}$ $(q-1)^{2}(q-2)/6$ $(q-1)(q^{2}-q+4)/6$ $q(q-1)^{2}/6$ $(q^{2}-1)(q-2)/6$
$A_{q-1}$ $(q-1)(q^{3}-q^{2}+14q-8)/6$ $(q-1)(q-2)(q^{2}+q+8)/6$ $q(q-1)(q^{2}-q+6)/6$ $(q^{2}-1)(q^{2}-2q+8)/6$
$A_{q}$ $3(q-1)(q^{3}-q+2)/2$ $(q-1)(3q^{2}+q+10)/2$ $(q-1)(3q^{2}+q+2)/2$ $(q^{2}-1)(3q-2)/2$
$A_{q+1}$ $(q-1)^{2}(3q^{2}+2q+8)/6$ $(q-1)(q-2)(3q^{2}+5q+8)/6$ $q(q-1)^{2}(3q+2)/6$ $(q-1)(3q^{3}-q^{2}-2q+8)/6$
$A_{q+2}$ $(q-1)^{3}(q+1)/3$ $(q^{2}-1)(q^{2}-2q+2)/3$ $q^{2}(q-1)^{2}/3$ $(q-1)(q^{3}-q^{2}-1)/3$
Condition $q\equiv\pm 1\pmod{3}$ $q\equiv-1\pmod{3}$ $q\equiv 1\pmod{3}$ $q\equiv-1\pmod{3}$
Remark 4.27.
Let $\mathbf{g}^{\mathrm{T}}_{q+2}=(0,0,1,0)\in\mathcal{M}_{2}$ in (4.1). Then we can directly obtain the corresponding result from [11, Theorem 15] using Theorem 4.26.
Example 4.28.
Let $q=8,h=1$, and take $\mathbf{g}^{\mathrm{T}}_{q+2}$ as $(0,1,0,0),(1,0,1,1),(1,\zeta,1,1)$ respectively, where $\zeta$ is a generator of $\mathbb{F}^{*}_{8}$. Using Magma [7], we obtain three $[10,4,6]$ NMDS codes with the following weight enumerators:
$A(z)=1+49z^{6}+644z^{7}+609z^{8}+1764z^{9}+1029z^{10},$
$A(z)=1+70z^{6}+560z^{7}+735z^{8}+1680z^{9}+1050z^{10},$
$A(z)=1+63z^{6}+588z^{7}+693z^{8}+1708z^{9}+1043z^{10},$
respectively, which coincides with the conclusion of Theorem 4.26.
Let $q=16,h=1$, and take $\mathbf{g}^{\mathrm{T}}_{q+2}$ as $(0,1,0,0),(1,\zeta^{3},1,1)$ respectively, where $\zeta$ is a generator of $\mathbb{F}^{*}_{16}$. Using Magma [7], we obtain two $[18,4,14]$ NMDS codes with the following weight enumerators:
$A(z)=1+525z^{14}+10140z^{15}+5445z^{16}+30300z^{17}+19125z^{108},$
$A(z)=1+600z^{14}+9840z^{15}+5895z^{16}+30000z^{17}+19200z^{108},$
respectively, which coincides with the conclusion of Theorem 4.26.
4.2 $q$ odd
In this subsection, we consider the case with odd $q$. Recall that if $q$ is odd every $(q+1)$-arc is a twisted cubic. Thus, we suppose that $(q+1)$-arc is the point-set
$\mathcal{S}_{1}=\{(1,x,x^{2},x^{3})\mid x\in\mathbb{F}_{q}\}\cup\{(0,0,0,1)\}.$
Similarly to the even case, let $K_{q}$ be the stabilizer of $\mathcal{S}_{1}$ in $\mathrm{PGL}(4,q)$. Then for $q>4$, $K_{q}$ is also isomorphic to $\mathrm{PGL}(2,q)$, and its elements are represented by the matrices
$M_{a,b,c,d}=\left(\begin{array}[]{cccc}a^{3}&3a^{2}c&3ac^{2}&c^{3}\\ a^{2}b&a^{2}d+2abc&bc^{2}+2acd&c^{2}d\\ ab^{2}&b^{2}c+2abd&ad^{2}+2bcd&cd^{2}\\ b^{2}&3b^{2}d&3bd^{2}&d^{3}\end{array}\right)\text{, }$
[…]
Lemma 4.29.
[17, Lemma 21.1.3][4, Theorem 2.2] Let $q\equiv 0\pmod{3}$. The group $K_{q}$ has five orbits on the points of $\mathrm{PG}(3,q)$:
  1. (1)
    $\mathcal{M}_{1}:=\mathcal{S}_{1}$ of size $q+1$;
  2. (2)
    $\mathcal{M}_{2}$ of size $q+1$, consisting of points on all osculating planes;
  3. (3)
    $\mathcal{M}_{3}$ of size $q^{2}-1$, consisting of points on exactly one osculating plane;
  4. (4)
    $\mathcal{M}_{4}$ of size $(q^{3}-q)/2$, consisting of points on a real chord;
  5. (5)
    $\mathcal{M}_{5}$ of size $(q^{3}-q)/2$, consisting of points on an imaginary chord.
Theorem 4.30.
Let $\mathscr{C}$ be the set of linear codes over $\mathbb{F}_{q}$ with generator matrices of the form (4.1). If $q\not\equiv 0\pmod{3}$, then there are two or three equivalence classes of NMDS codes in $\mathscr{C}$, depending on whether $q\equiv 1\pmod{3}$ or $q\equiv-1\pmod{3}$. Moreover, the weight distributions of these codes are the same as those in Theorem 4.26. If $q\equiv 0\pmod{3}$, then there are three equivalence classes of NMDS codes in $\mathscr{C}$, and their weight distributions are given in Table 3.
Proof.
Note that the number of planes containing exactly three points of $\mathcal{S}_{1}$ through a point of $\mathcal{M}_{i}$ can be found in [4, Table 1, 2]. To prove the theorem, it suffices to show that in the case $q\equiv 0\pmod{3}$, the linear codes corresponding to $\mathbf{g}_{q+2}$ in the $K_{q}$-orbits $\mathcal{M}_{2}$ and $\mathcal{M}_{5}$ are not monomially equivalent. The remaining proofs follow similarly to those in Theorem 4.26. Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be linear codes with generator matrices $G_{1}=(\mathbf{g}_{1},\ldots,\mathbf{g}_{q+1},\mathbf{e})$ and $G_{2}=(\mathbf{g}_{1},\ldots,\mathbf{g}_{q+1},\mathbf{f})$ respectively, where $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{q+1}\}$ forms the $(q+1)$-arc $\mathcal{S}_{1}$ and $\mathbf{e}\in\mathcal{M}_{2},\mathbf{f}\in\mathcal{M}_{5}$. Suppose that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are monomially equivalent. Then there is a map $\theta\in\mathrm{PGL}(4,q)$ such that the sets of projective points $\{\mathbf{g}^{\theta}_{1},\ldots,\mathbf{g}^{\theta}_{q+1},\mathbf{e}^{\theta}\}=\{\mathbf{g}_{1},\ldots,\mathbf{g}_{q+1},\mathbf{f}\}$. Obviously, $\theta\notin K_{q}$ since $\mathbf{e}$ and $\mathbf{f}$ belongs to different $K_{q}$-orbits. Therefore, we must have $\mathbf{e}^{\theta}=\mathbf{g}_{i}$ for some $i\in\{1,\ldots,q+1\}$. By Lemma 4.29, the point $\mathbf{e}$ lies on all osculating planes and so does $\mathbf{e}^{\theta}=\mathbf{g}_{i}$, which is impossible. Thus $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are not monomially equivalent. This completes the proof. ∎
Table 3: The weight distributions of NMDS codes in Theorem 4.30 when $\mathbf{g}_{q+2}\in\mathcal{M}_{i}$ and $q\equiv 0\pmod{3}$.
$\mathcal{M}_{2}$ $\mathcal{M}_{3}$ $\mathcal{M}_{5}$
$A_{q-2}$ $q(q-1)^{2}/6$ q(q1)(q3)/6q(q-1)(q-3)/6 $q(q-1)^{2}/6$
$A_{q-1}$ $q(q-1)(q^{2}-q+6)/6$ $q(q-1)(q^{2}-q+14)/6$ $q(q-1)(q^{2}-q+6)/6$
$A_{q}$ $(q-1)(3q^{2}+q+2)/2$ $(q-1)(3q^{2}-3q+2)/2$ $(q-1)(3q^{2}+q+2)/2$
$A_{q+1}$ $q(q-1)^{2}(3q+2)/6$ $q(q-1)(3q^{2}-q+6)/6$ $q(q-1)^{2}(3q+2)/6$
$A_{q+2}$ $q^{2}(q-1)^{2}/3$ $q(q-1)(q^{2}-q-1)/3$ $q^{2}(q-1)^{2}/3$
4.3 NMDS codes with parameters [q+3,4,q1][q+3,4,q-1]
Let $q=p^{m}$ be a prime power and $h$ be a positive integer. Define $4\times(q+3)$ matrix $G_{3}$ by
$G_{3}=(\mathbf{g}_{1},\ldots,\mathbf{g}_{q+1},\mathbf{g}_{q+2},\mathbf{g}_{q+3})=\begin{pmatrix}1&1&\cdots&1&1&0&a_{1}&a_{2}\\ \alpha_{1}&\alpha_{2}&\cdots&\alpha_{q-1}&0&0&b_{1}&b_{2}\\ \alpha_{1}^{2^{h}}&\alpha_{2}^{2^{h}}&\cdots&\alpha_{q-1}^{2^{h}}&0&0&c_{1}&c_{2}\\ \alpha_{1}^{2^{h}+1}&\alpha_{2}^{2^{h}+1}&\cdots&\alpha_{q-1}^{2^{h}+1}&0&1&d_{1}&d_{2}\end{pmatrix}$, (4.2)
where $\mathbb{F}^{*}_{q}=\mathbb{F}_{q}\setminus\{0\}=\{\alpha_{1},\dots,\alpha_{q-1}\}$ and $\mathbf{g}_{q+2},\mathbf{g}_{q+3}\neq\mathbf{0}$. By selecting specific $\mathbf{g}_{q+2}$ and $\mathbf{g}_{q+3}$, we can construct certain NMDS codes. For example, let $p=2$ and $h$ be a positive integer satisfying $\gcd(m,h)=1$, where $m\geq 3$ is odd. Take $\mathbf{g}_{q+2}=(0,0,1,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ and $\mathbf{g}_{q+3}=(0,1,0,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ or $\mathbf{g}_{q+3}=(0,1,1,0)^{\mathrm{T}}\in\mathcal{M}_{3}$, where $\mathcal{M}_{2}$ and $\mathcal{M}_{3}$ are the point-orbits of $K_{h}$ defined in Lemma 4.24. Then the linear code generated by $G_{3}$ is a [q+3,4,q1][q+3,4,q-1] NMDS code (see [11, Theorems 18 and 19]). A simpler proof of this result can be provided via Theorem 4.26.
Theorem 4.31.
Let $q=2^{m}$ and $h$ be a positive integer satisfying $\gcd(m,h)=1$. Take $\mathbf{g}_{q+2}=(0,0,1,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ and $\mathbf{g}_{q+3}=(0,1,0,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ in (4.2). Let $\mathcal{C}_{3}$ be the linear code generated by $G_{3}$. Then $\mathcal{C}_{3}$ is a [q+3,4,q1][q+3,4,q-1] NMDS code with $A_{q-1}=(q-1)^{2}(q-2)/3$ if and only if $m\geq 3$ is odd.
Proof.
First, we can verify that any three columns of $G_{3}$ are linearly independent, i.e. $\mathcal{C}_{3}$ satisfy condition (1). Then, for any nonzero vector $u=(u_{1},u_{2},u_{3},u_{4})\in\mathbb{F}^{4}_{q}$, we calculate $|\mathcal{H}_{u}\cap\mathcal{S}_{G_{3}}|$, where $\mathcal{S}_{G_{3}}$ denotes the set of points $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{q+1},\mathbf{g}_{q+2},\mathbf{g}_{q+3}\}$. Suppose that $\mathbf{g}_{q+2}\in\mathcal{H}_{u}$ and $\mathbf{g}_{q+3}\notin\mathcal{H}_{u}$. By the proof of Theorem 4.26, we have $|\mathcal{H}_{u}\cap\mathcal{S}_{G_{3}}|\leq 4$ and the number of nonzero vectors $u$ satisfying this condition is $(q-1)^{2}(q-2)/6$ from Table 2. The case where $\mathbf{g}_{q+2}\notin\mathcal{H}_{u}$ and $\mathbf{g}_{q+3}\in\mathcal{H}_{u}$ also yields the same results.
Finally, suppose that $\mathbf{g}_{q+2},\mathbf{g}_{q+3}\in\mathcal{H}_{u}$. Then $u_{2}=u_{3}=0$. If $(1,x,x^{2^{h}},x^{2^{h}+1})\in\mathcal{H}_{u}$, then we have $u_{1}+u_{4}x^{2^{h}+1}=0$. If $m$ is odd, then $\gcd(2^{h}+1,q-1)=\gcd(2^{h}+1,2^{m}-1)=1$ since $\gcd(h,m)=1$. By Lemma 2.7, the equation $u_{1}+u_{4}x^{2^{h}+1}=0$ has exactly one solution, implying $|\mathcal{H}_{u}\cap\mathcal{S}_{G_{3}}|=3$. Therefore, $|\mathcal{H}_{u}\cap\mathcal{S}_{G_{3}}|\leq 4$ holds for any nonzero vector $u\in\mathbb{F}^{4}_{q}$, and the number of vectors $u\in\mathbb{F}^{4}_{q}$ such that $|\mathcal{H}_{u}\cap\mathcal{S}_{G_{3}}|=4$ is $(q-1)^{2}(q-2)/6+(q-1)^{2}(q-2)/6=(q-1)^{2}(q-2)/3$. Thus, $\mathcal{C}_{3}$ is a [q+3,4,q1][q+3,4,q-1] NMDS code with $A_{q-1}=(q-1)^{2}(q-2)/3$. If $m$ is even, then $\gcd(2^{h}+1,q-1)=\gcd(2^{h}+1,2^{m}-1)=2^{\gcd(h,m)}+1=3$. By Lemma 2.7, the equation $u_{1}+u_{4}x^{2^{h}+1}=0$ has exactly three solutions, leading to $|\mathcal{H}_{u}\cap\mathcal{S}_{G_{3}}|=5$. Hence, $\mathcal{C}_{3}$ is not an NMDS code. This completes the proof. ∎
Theorem 4.32.
Let $q=2^{m}$ and $h$ be a positive integer satisfying $\gcd(m,h)=1$. Take $\mathbf{g}_{q+2}=(0,0,1,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ and $\mathbf{g}_{q+3}=(0,1,1,0)^{\mathrm{T}}\in\mathcal{M}_{3}$ in (4.2). Let $\mathcal{C}_{4}$ be the linear code generated by $G_{3}$. Then $\mathcal{C}_{4}$ is a [q+3,4,q1][q+3,4,q-1] NMDS code with $A_{q-1}=(q-1)(q^{2}-2q+3)/3$ if and only if $m\geq 3$ is odd.
Proof.
The proof follows the approach of Theorem 4.31. The value of $A_{q-1}$ is obtained by summing the entries in the second column $(q-1)^{2}(q-2)/6$ and the third column $(q-1)(q^{2}-q+4)/6$ of the second row in Table 2. ∎
Let $h=1$ and $\gcd(3,q-1)=1$. Taking $\mathbf{g}_{q+2}=(0,0,1,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ and $\mathbf{g}_{q+3}=(0,1,0,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ in (4.2), the linear code generated by $G_{3}$ is a [q+3,4,q1][q+3,4,q-1] NMDS code (see [41, Theorem 9]). Similarly to the proof of Theorem 4.31, we further have the following Theorem from Tables 2 and 3.
Theorem 4.33.
Let $h=1$ and $q=p^{m}$. Take $\mathbf{g}_{q+2}=(0,0,1,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ and $\mathbf{g}_{q+3}=(0,1,1,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ in (4.2). Let $\mathcal{C}_{5}$ be the linear code generated by $G_{3}$. Then $\mathcal{C}_{5}$ is a [q+3,4,q1][q+3,4,q-1] NMDS code if and only if $\gcd(3,q-1)=1$. Moreover, $A_{q-1}=q(q-1)^{2}/3$ if $p=3$ and $A_{q-1}=(q-1)^{2}(q-2)/3$ if $p\neq 3$.
When $q=2^{m}$ with $m$ odd, Theorems 4.31 and 4.33 show that $\mathcal{C}_{3}$ and $\mathcal{C}_{5}$ have identical weight enumerators by Lemma 2.2. However, as noted in [11, Remark 22], $\mathcal{C}_{3}$ and $\mathcal{C}_{5}$ are not monomially equivalent [q+3,4,q1][q+3,4,q-1] NMDS codes for $q=2^{3},2^{5},2^{7}$ and some possible values of $h$. Finally, we have the following proposition.
Proposition 4.34.
If $q=2^{m}$ with $m\geq 7$ and $h\geq 2$, then $\mathcal{C}_{3}$ and $\mathcal{C}_{5}$ are not monomially equivalent.
Proof.
Suppose that $\mathcal{C}_{3}$ and $\mathcal{C}_{5}$ are monomially equivalent. Let $\mathcal{S}_{h}$ and $\mathcal{S}_{1}$ be the $(q+1)$-arcs corresponding to the first $q+1$ columns of generator matrices of $\mathcal{C}_{3}$ and $\mathcal{C}_{5}$, respectively. By analogy to Lemma 3.10, there exists a map $\theta\in\mathrm{PGL}(4,q)$ such that $\mathcal{S}^{\theta}_{h}\cup\{\mathbf{g}^{\theta}_{q+2},\mathbf{g}^{\theta}_{q+3}\}=\mathcal{S}_{1}\cup\{\mathbf{g}_{q+2},\mathbf{g}_{q+3}\}$. This implies that $|\mathcal{S}^{\theta}_{h}\cap\mathcal{S}_{1}|\geq q-3>q-\sqrt{q}/2+9/4$. By Lemma 2.8, it follows that $\mathcal{S}^{\theta}_{h}=\mathcal{S}_{1}$, i.e., $\mathcal{S}_{h}$ is projectively equivalent to $\mathcal{S}_{1}$, which is a contradiction. This completes the proof. ∎
5 Conclusion
In this paper, we investigated the monomial equivalence and construction of NMDS codes with dimensions 3 and 4 using group theory and geometric methods. In dimension 3, we proposed a criterion for the monomial equivalence of NMDS codes derived from hyperovals in $\mathrm{PG}(2,q)$. By analyzing the actions of the homography stabilizers of distinct hyperovals on the points not in hyperovals, we determined the equivalence classes of NMDS codes derived from augmenting hyperovals with a single point. In dimension 4, we also determined the equivalence classes of NMDS codes derived from arcs in $\mathrm{PG}(3,q)$ with additional points, which included new constructions of NMDS codes of length $q+2$. By analyzing the orbits of points and planes of $\mathrm{PG}(3,q)$ under the stabilizers of arcs, we provided a unified geometric perspective for prior constructions of NMDS codes (e.g., [11, 41]) and simplified the proofs through geometric invariants instead of algebraic computations. Finally, while the stabilizer-based approach is highly effective for codes derived from arcs or hyperovals, its applicability to longer, less structured NMDS codes remains a challenging and interesting open problem.
6 Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant No. 12501467 and the National Key Research and Development Program of China under Grant No. 2022YFA1004900.